OpenStudy (anonymous):

please help!!! What is the solution to the equation 2^1 - 2x = 3 ?

OpenStudy (luigi0210):

Well x^1=x So that means 2^1=2

OpenStudy (zzr0ck3r):

2^1=2 so 2-2x=3 can you solve this?

OpenStudy (luigi0210):

@shannondidier follow off of what zz did

OpenStudy (anonymous):

i got -2, but my answer choices are x ≈ -0.292 x ≈ -0.185 x ≈ 0.185 x ≈ 0.292

OpenStudy (luigi0210):

Is the x an exponent too?

OpenStudy (anonymous):

the 1-2x is an exponent

OpenStudy (luigi0210):

@zzr0ck3r

OpenStudy (anonymous):

I believe the original problem is \[2^{1-2x}=3\]

OpenStudy (anonymous):

it is

OpenStudy (zzr0ck3r):

2^(1-2x) = 3 1-2x=log_2(3) 1-log_2(3)=2x x=(1-log_2(3))/2

OpenStudy (zzr0ck3r):

about -0.292

OpenStudy (anonymous):

thank you so much, i understand now! :)

OpenStudy (zzr0ck3r):

good deal